LIST OF BAD FIBERS/TILES FOR ANL009
 

QC 6 VERSION 4.1

First Scan: 9-Oct-00

PMT

LAYER

CELL

TILE

SLOT

%DEV

SOLUTION

             

3

10

D4-O

17

17

-27.8

Reseated Profile

8

3

A12-E

9: 9

26

MISSING

On side of tile - added shim, reseated profile

16

5

B12-E

13

46

-72.9

No visible reason, Replaced fiber

16

7

B12-E

13

46

-72.8

17

9

D5-O

9

26

-26.1

Reseated Profile

17

9

D5-O

35

52

-31.1

Fiber detached from profile - reglued

17

11

D5-O

35

52

-28.1

17

11

D5-O

65

82*

-39.8

On side of tile - added shim, reseated profile

18

9

D5-E

65:65

82*

MISSING

On side of tile - added shim, reseated profile

22

2

A14-E

22

73

-26.8

Reseated Profile

29

1

A15-O

21

100

-27.6

15 mm of glue on end of fiber - not replaced

29

3

A15-O

21

100

-27.4

37

9

D6-O

48

130*

-46.3

On side of tile - reseated profile

37

10

D6-O

63

145

-77.5

On side of tile - reseated profile

37

10

D6-O

64

146

-64.7

On side of tile - reseated profile

37

10

D6-O

70

152

-74.0

On side of tile - reseated profile

38

8

D6-E

42

124

-25.7

Profiles and fibers look OK. Checked polished end of aspirin tube and found some paint flakes. Cleaned it.

38

9

D6-E

41

123

-27.9

38

9

D6-E

53

135

-32.8

38

9

D6-E

54

136

-38.3

38

9

D6-E

58

140

-27.5

38

9

D6-E

59

141

-32.1

38

9

D6-E

61

143

-36.0

38

11

D6-E

13

95

-48.0

On side of tile - reseated profile

38

11

D6-E

48

130*

-38.9

On side of tile - added shim, reseated profile

38

11

D6-E

53

135

-27.2

See note for these fibers in layer 9.

38

11

D6-E

54

136

-30.1

38

11

D6-E

59

141

-26.1

38

11

D6-E

61

143

-28.9

42

2

A16-E

34

143

-37.0

These 4 fibers had been glued to the profile with too much glue. They were replaced and repolished

42

2

A16-E

37

146

-52.0

44

5

B15-E

19

141

-30.3

44

5

B15-E

20

142*

-33.9

44

7

B15-E

19

141

-28.0

44

7

B15-E

20

142*

-32.5

 

Second LED Scan: 16-Oct-00

 

PMT

LAYER

CELL

TILE

SLOT

%DEV

SOLUTION

4

10

D4-E

17

34

MISSING

These are not real - caused by large pedestal shifts on PMT 4

4

10

D4-E

18

36*

MISSING

4

10

D4-E

19

38

-50.7

17

9

D5-O

9

51

-25.0

 

17

9

D5-O

35

103

-32.2

Fiber that was re-glued

17

11

D5-O

35

103

-31.4

29

3

A15-O

21

199

-25.8

See first scan above

38

9

D6-E

41

245

-25.4

These fibers on neighboing tiles appear to have been reglued to profiles before those profiles were installed in this module.

At ANL, we replaced this set of tiles with brighter tiles from production batch 3 and were able to improve the uniformity for most of this region

38

9

D6-E

53

269

-30.0

38

9

D6-E

54

271

-34.6

38

9

D6-E

59

281

-26.5

38

9

D6-E

61

285

-33.2

38

11

D6-E

53

269

-28.1

38

11

D6-E

54

271

-29.6

38

11

D6-E

58

279

-27.0

38

11

D6-E

59

281

-25.1

38

11

D6-E

61

285

-29.9

42

1

A16-E

13

243

-29.0

New on this list

42

2

A16-E

23

264

-26.4

New on this list

42

3

A16-E

13

243

-25.2

New on this list